Given angle 1 is complement of angle 2 angle 2 is congruent to angle 3 prove angle 1 is a complement of angle 3

Answer by math_tutor2020(3835)


Angles 1 and 2 are complementary
This means they add to 90
(angle1)+(angle2) = 90

Because angles 2 and 3 are congruent, this means they are the same. We can replace "angle2" with "angle3" through the substitution rule.

We go from this
(angle1)+(angle2) = 90
to this
(angle1)+(angle3) = 90
which shows that angles 1 and 3 are complementary.